3. The length of the minute hand of a clock is 14 cm. Find the area swept by the minute hand in 5 minutes.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Length of minute hand = radius of the clock (circle) ∴ Radius (r) of the circle = 14 cm (given) Angle swept by minute hand in 60 minutes = 360° So, the angle swept by the minute hand in 5 minutes = 360° × 5/60 = 30° We know, Area of a sector = (θ/360°) × πr2 Now, area of the sector making an angle of 30° = (30°/360°) × πr2 cm2 = (1/12) × π142 = (49/3)×(22/7) cm2 = 154/3 cm2